OpenStudy (anonymous):

how to simplify fractions help plz any fraction make up ur own and plz try an use some negatives plz do three one like atachment

OpenStudy (anonymous):

thanks alot i dont get fractions at all your all a big help bye bye

OpenStudy (anonymous):

\[\frac{ -103 }{25 }=\frac{ -103 }{25 }*\frac{ 4 }{ 4 }=\frac{ -412 }{ 100 }=-4.12\] \[-\frac{ 103 }{ 25 }=-4\frac{ 3 }{ 25 }\] here i have given the answer in decimals & a proper fraction \[\frac{ 10 }{-9 }=-1\frac{ 1 }{ 9 }\] &so on

OpenStudy (anonymous):

i dont get it sorry 6 grader here

OpenStudy (anonymous):

lets say we have the fraction \[\frac{128}{24}\]

OpenStudy (anonymous):

we want to look at the factors of 128 and the factors of 24 an obvious factor would be 2, because both numbers are even and thus divisible by 2 so we pull out a 2 from both numbers \[\frac{128}{24}= \frac{ 2*64}{2*12}\] now since 2/2 = 1 \[\frac{ 2*64}{2*12}=1* \frac{64}{12} = \frac{64}{12}\]

OpenStudy (anonymous):

then you repeat the process looking at the factors of 64 and 12,